OpenStudy (anonymous):

prove that (2n)!/n!^2 is an even integer for n>0

OpenStudy (anonymous):

Use induction. First show it's true for n = 1 Then show that if (2n)!/n!^2 is even, then (2(n+1))!/(n+1)!^2 must be even

OpenStudy (anonymous):

how about using binomial coefficients?

OpenStudy (anonymous):

how u would use induction to show it must be true (2n+2)!/(n+1)!^2

myininaya (myininaya):

Try to rewrite (2k+2)! and (k+1)! as blah*blah*(2k)! and (k+1)*k! respectively I would let you try to figure the blah blah out.

OpenStudy (anonymous):

((2k+2)*(2+1) *2k! )/(k+1)*k!

OpenStudy (anonymous):

((2k+2)*(2k+1) *2k! )/(k+1)*k!

OpenStudy (anonymous):

and what does it give us?

myininaya (myininaya):

by induct that one part is even.

myininaya (myininaya):

something*even=still an even number

myininaya (myininaya):

also you forgot the square part.

myininaya (myininaya):

\[\frac{(2k+2)(2k+1) \cdot (2k)!}{(k+1)^2 (k!)^2}=\frac{(2k+2)(2k+1)}{(k+1)^2} \cdot \frac{(2k)!}{(k!)^2} \] But we really didn't even need that induction step. Like we could have looked at that other factor that contains the factor 2, but whateves...either way.

OpenStudy (anonymous):

im kind of confused

OpenStudy (anonymous):

how would the inductive step look because I'm still confused? So assume that the equation is true then we have to proof that (2(k+1))!/(k+1)!^2 is true?

myininaya (myininaya):

Yes. (2(k+1))!=(2k+2)!=(2k+2)*(2k+1)*(2k)! [(k+1)!]^2=[(k+1)*k!]^2=(k+1)^2*[k!]^2

OpenStudy (anonymous):

but then we still don't know that even is multiplied by the integer

myininaya (myininaya):

You can play with the expression if you want to convince yourself it is an integer.

myininaya (myininaya):

Re-write the top and the bottom so you see the bottom stuff cancel out.

OpenStudy (anonymous):

sorry for so many questions, but i don't see how. we have to leave \[\frac{(2k)!}{(k!)^2}\] alone because thats our induction so we are left with \[\frac{ (2k+2)(2k+1) }{(k+1)(k+1) }\] and the only way i can think of simplifying that is \[4-\frac{2}{k+1}\] but we are still left with k+1